# Dilaton Effective Field Theory across the Conformal Edge

### v1

@@ I Introduction
In four dimensions, asymptotically free gauge theories are realized in different zero-temperature phases, depending on the type and number $N_{f}$ of light matter fields included. In the absence of matter fields, Yang-Mills theories confine, and the presence of few fermions does not alter qualitatively this feature, except for introducing continuous global symmetries and their spontaneous breaking. Conversely, for a large enough number of fermions, long distance physics is governed by the appearance of an infrared fixed point, with its associated conformal invariance. This can be shown perturbatively, when $N_{f}$ is close to its upper bound insuring asymptotic freedom [Banks:1981nn]. The conformal behavior is expected to persist over a finite range of $N_{f}>N_{f\,c}$, with the lower bound defining the edge of the conformal window. The determination of $N_{f\,c}$, and the characterization of the theory near it, is an interesting strong-coupling problem.
@@
Lattice field theory is the natural instrument to study the problem, and in recent years it has led to significant advances—see the review in Ref. [Rummukainen:2022ekh] and references therein, as well as the more recent Refs. [Chung:2023mgr, Bennett:2023gbe, Nogradi:2023wnf, Bennett:2024tex, Athenodorou:2024rba]. These studies explore a finite range of fermion masses that suppress infrared sensitivity to the lattice volume, and yield a discrete spectrum of bound states, as expected in confining theories. Some of these studies have reported evidence for the existence of a relatively light, flavor-singlet particle, together with a set of similarly light pseudoscalar particles, separated in mass from heavier bound states [Appelquist:2007hu, Deuzeman:2008sc, Fodor:2009wk, LatKMI:2016xxi, Appelquist:2016viq, LatticeStrongDynamics:2018hun, Kotov:2021mgp, LatticeStrongDynamicsLSD:2021gmp, Hasenfratz:2022qan, LatticeStrongDynamics:2023bqp, LSD:2023uzj, LatKMI:2025kti, Fodor:2012ty, Fodor:2015vwa, Fodor:2016pls, Fodor:2017nlp, Fodor:2019vmw, Fodor:2020niv, Athenodorou:2014eua, Athenodorou:2016ndx, Athenodorou:2021wom, Athenodorou:2024rba].
@@
In a gauge theory outside of the conformal window, the pseudoscalars are naturally interpreted as pseudo-Nambu-Goldstone Bosons (pNGBs), associated with the spontaneous breaking of an approximate internal global symmetry acting on the matter fields. With the light flavor-singlet interpreted as a dilaton associated with the spontaneous breaking of approximate dilatation symmetry, the behavior of a gauge theory taken to be just below the edge of the window has been successfully described employing dilaton effective field theory (dEFT) [Matsuzaki:2013eva, Golterman:2016lsd, Kasai:2016ifi, Hansen:2016fri, Golterman:2016cdd, Appelquist:2017wcg, Appelquist:2017vyy, Cata:2018wzl, Golterman:2018mfm, Cata:2019edh, Appelquist:2019lgk, Golterman:2020tdq, Golterman:2020utm]. See also the review in Ref. [Appelquist:2022mjb] and references therein, including the precursors in Refs. [Migdal:1982jp, Coleman:1985rnk], the early study in the context of electroweak symmetry breaking [Goldberger:2007zk], as well as recent phenomenological applications in Refs. [Appelquist:2020bqj, Appelquist:2022qgl, Cacciapaglia:2023kat, Appelquist:2024koa, Bruggisser:2022ofg, Alonso:2025ksv, Wu:2025hfp], and theoretical developments in Refs. [Zwicky:2023krx, Faedo:2024zib, Elander:2025fpk, Cresswell-Hogg:2025kvr, Stegeman:2025sca, Stegeman:2025tdl].
@@
In this paper, we expand on this idea, noting that the same dEFT can also be used just inside the conformal window, in the presence of a finite fermion mass. The dilaton and pNGBs continue to exist there as relatively light states. Therefore, dEFT can be employed to explore near-edge gauge theories without an inside-versus-outside bias, and hence used as a diagnostic tool to discriminate between confining and conformal behavior. We test this framework by applying it to two gauge theories in proximity to the lower edge of the conformal window. We start with the well studied $SU(3)$ gauge theory with $N_{f}=8$ Dirac fermions in the fundamental representation [LatticeStrongDynamics:2023bqp, LSD:2023uzj], finding, as assumed in Refs. [Matsuzaki:2013eva, Golterman:2016lsd, Kasai:2016ifi, Hansen:2016fri, Golterman:2016cdd, Appelquist:2017wcg, Appelquist:2017vyy, Cata:2018wzl, Golterman:2018mfm, Cata:2019edh, Appelquist:2019lgk, Golterman:2020tdq, Golterman:2020utm], that it is outside the conformal window. We then analyze the $SU(2)$ gauge theory coupled to $N_{f}=1$ Dirac fermion transforming in the adjoint representation [Athenodorou:2021wom, Athenodorou:2024rba], finding evidence that it is inside.
@@
In Section II, we discuss the general framework. We then implement it to fit the lattice data for the above two gauge theories in Section III. In Section IV, we summarize our results and discuss future directions. We relegate to the Appendix some technical details about our fitting procedure, and tables reproducing the data we take from Refs. [LatticeStrongDynamics:2023bqp, SU(3)-data_release] and [Athenodorou:2021wom, SU(2)-1-data_release, SU(2)-2-data_release].
@@ II Dilaton Effective Field Theory
The matrix field $\Sigma$ describes the pNGBs. The factor $\chi^{2}$ multiplying their kinetic term acts as a conformal compensator, and $P$ is a positive dimensionless parameter to be fit to lattice data. The final term describes explicit breaking of the internal global symmetry arising from the finite mass $m$ assigned to the fermions in the underlying gauge theory. The $\chi$ dependence of the terms in Eq. (II) follows from a spurion analysis, explained in Refs. [Appelquist:2019lgk, Appelquist:2022mjb]. Heavier states and subleading interactions are neglected.
@@
We note in passing that our previous studies [Appelquist:2019lgk, Appelquist:2022mjb] of the $SU(3)$ gauge theory with $N_{f}=8$ fermions in the fundamental representation were conducted with a constraint on the parameters of the potential $V(\chi)$ such that it has a symmetry breaking minimum, insuring that the spontaneous breaking of approximate conformal symmetry persists in the limit $R=0$. Here we proceed more generally, allowing lattice fits to determine whether or not $V(\chi)$ has a symmetry-breaking minimum, that is whether the theory is just outside or just inside the conformal window.
@@ III Fits to Lattice Data
Figure 3: Plot of $a^{3}M^{2}_{\pi}F_{\pi}$ versus $aF_{\pi}$, in the $SU(3)$ theory with $N_{f}=8$ fermions in the fundamental representation [LatticeStrongDynamics:2023bqp, SU(3)-data_release]. The red curve is based on the central values of Table 1, with the five data points denoted by the black discs. The blue shaded region represents the uncertainty in the fitting curve, equivalent to the $1\sigma$ confidence level. The red curve reaches the horizontal axis at a nonzero value of $aF_{\pi}$, corresponding to the limit $m\rightarrow 0$ in a confining gauge theory. For larger values of $F_{\pi}\propto F_{d}$, we have $M_{\pi}^{2}F_{\pi}\propto V^{\prime}(F_{d})$. The curve is extrapolated downward, where it becomes negative, by continuing Eq. (10) to unphysical values of $F_{\pi}$, demonstrating that the region near the origin is unstable, as shown in the inset.
Table 1: Central values in the dEFT fit for the $SU(3)$ gauge theory with $N_{f}=8$ Dirac fermions in the fundamental representation. Lattice data for $M_{\pi}$, $M_{d}$, and $F_{\pi}$ is incorporated into this fit, for 5 different values of the underlying fermion mass $m$, corresponding to 15 data points. The lattice measurements are taken from Refs. [LatticeStrongDynamics:2023bqp, SU(3)-data_release], and reproduced in Table 3. Dimensional quantities are presented in units of the lattice spacing, $a$.
@@ III.1 $SU(3)$ Gauge Theory with $N_{f}=8$ Fermions in the Fundamental Representation
The lattice data obtained from Refs. [LatticeStrongDynamics:2023bqp, SU(3)-data_release] is summarized in Table 3 of Appendix A. The masses and decay constants of the $63$ pNGBs and the masses of the dilaton are listed there for $5$ values of the underlying fermion mass $m$. As noted in the Appendix, correlations are neglected. The choice of lattice coupling $\beta$ and the range of fermion mass $m$ employed in the lattice study lead to masses $M_{\pi}$ and $M_{d}$ well below the inverse lattice spacing, and also below the masses of heavier states, which are not included in the dEFT analysis.
Results for $A$, $B$, and $\Delta$ are shown in Figs. 1 and 2 in the form of $\Delta\chi^{2}$ contours. The favored upper left quadrants together correspond to values $\Delta<4$, $A>0$, and $B<0$. The central values of all $6$ fit parameters are shown in Table 1. Those for $y$ and $P$ are consistent with earlier studies. The central values of $\Delta$ (<4)<4), AA (positive), and $B$ (negative) imply that the theory lies just outside the conformal window, with $V(\chi)$ having a symmetry breaking minimum, as assumed in our earlier studies [Appelquist:2022mjb].
@@ III.2 $SU(2)$ Gauge Theory with $N_{f}=1$ Fermion in the Adjoint Representation
Figure 6: Plot of $a^{3}M^{2}_{\pi}F_{\pi}$ versus $aF_{\pi}\propto aF_{d}$, in the $SU(2)$ theory with $N_{f}=1$ fermion in the adjoint representation  [Athenodorou:2021wom, SU(2)-1-data_release, SU(2)-2-data_release], restricting attention to ensembles with $\beta=2.1$. The red curve is based on the central values of Table 2, using only the three lightest ensembles. The data points are depicted by the black discs. The red curve reaches the horizontal line at $F_{\pi}=0$, as shown in the inset, corresponding to the limit $m\rightarrow 0$ in an infrared conformal theory. Throughout the range, $a^{3}M^{2}_{\pi}F_{\pi}\propto V^{\prime}(F_{d})$. The blue shaded region represents the uncertainty in the fitting curve, equivalent to the $1\sigma$ confidence level.
Table 2: Central values for the dEFT fit for the $SU(2)$ gauge theory with $N_{f}=1$ Dirac fermion in the adjoint representation. The lattice data is taken from Refs. [Athenodorou:2021wom, SU(2)-1-data_release, SU(2)-2-data_release], and reproduced in Table 4, and has been restricted to the three ensembles with smallest fermion mass and $\beta=2.1$. Lattice data for $M_{\pi}$, $M_{d}$, $F_{\pi}$, and $m_{\rm PCAC}$ has been incorporated into this fit, corresponding to 9 data points. Dimensional quantities are presented in units of the lattice spacing, $a$.
@@
We again employ Eqs. (9)-(11) to fit the lattice data [Athenodorou:2021wom, Athenodorou:2024rba, SU(2)-1-data_release, SU(2)-2-data_release]. In this case, with the spontaneous breaking pattern $SU(2)\rightarrow SO(2)$, there are $2$ pNGBs. We list the measurements performed with Wilson fermions and coupling $\beta=2.1$, for the largest available volumes, in Table 4 of Appendix A. We identify the fermion mass $m_{\rm PCAC}$, defined following Ref. [DelDebbio:2007wk], with the mass parameter $m$ in our fits. We include in our analysis only the three ensembles with the lightest mass, $aM_{\pi,d}<0.7$. We tested the stability of the fits by including one or more heavier ensembles. [^2] As discussed in Ref. [Athenodorou:2024rba], ensembles with $\beta>2.1$ yield a systematic decrease in the anomalous dimension of the fermion condensate, indicating to us that in those ensembles the Wilson fermion mass does not reach low enough values to assess the existence of an IR fixed point.
Results for $A$, $B$, and $\Delta$ are shown in Figs. 4 and 5 in the form of $\Delta\chi^{2}$ contours, with central values listed in Table 2. Among the central values, those for $y$ and $P$ are similar to those of Table 1. In particular, $y\approx 2$, as expected near the edge of the conformal window [Cohen:1988sq]. We again have $A>0$ and $B<0$, but now $\Delta>4$. Thus $V(\chi)$ turns down at large $\chi$, but the turn-down sets in well beyond the range of applicability of the dEFT.
@@ IV Summary and Discussion
We employed dilaton effective field theory (dEFT) as a diagnostic tool to discriminate between infrared-conformal and near-conformal, confining gauge theories. Our approach draws on spectroscopic lattice measurements performed with finite mass for the elementary fermions, in a range of parameter space in which both the pNGBs and the dilaton are clearly lighter than other states. This feature is expected to emerge naturally in theories near the lower edge of the conformal window, for small enough fermion masses. We took the dEFT to apply smoothly through the transition between confinement and conformality.
We tested this approach with two lattice studies meeting these requirements. For the $SU(3)$ gauge theory coupled to $N_{f}=8$ Dirac fermions in the fundamental representation, we used measurements obtained with staggered fermions, restricted to the largest available lattice couplings and volumes [LatticeStrongDynamics:2023bqp, LSD:2023uzj], to infer that the theory is outside the conformal window (near-conformal but confining in the $m\rightarrow 0$ limit). For the $SU(2)$ theory with $N_{f}=1$ fermions in the adjoint representation, we used data obtained with Wilson fermions [Athenodorou:2021wom, Athenodorou:2024rba], by restricting attention to ensembles in which the relevant masses are significantly smaller than the lattice cutoff, to minimize lattice artifacts. Our analysis indicates that this theory is inside the conformal window (infrared conformal in the $m\rightarrow 0$ limit).
@@
Our results should be taken as preliminary, as indicated by the large uncertainty in some of the output—in particular the scaling dimension $\Delta$. Future precision lattice measurements, particularly for the $SU(2)$ theory, could overcome these limitations, by exploring lattice parameters with smaller fermion masses and less contamination from lattice artifacts. Our framework can be applied to other gauge theories. It would be interesting, for example, to perform our analysis on lattice measurements in the $SU(3)$ theory coupled to sextet fermions [Fodor:2012ty, Fodor:2015vwa, Fodor:2016pls, Fodor:2017nlp, Fodor:2019vmw, Fodor:2020niv], or theories in other gauge groups, such as $Sp(4)$ [Bennett:2023gbe, Bennett:2024tex].

### v2

@@ I Introduction
In four dimensions, asymptotically free gauge theories are realized in different zero-temperature phases, depending on the type and number $N_{f}$ of light matter fields included. In the absence of matter fields, Yang-Mills theories confine, and the presence of few fermions does not alter qualitatively this feature, except for introducing continuous global symmetries and their spontaneous breaking. Conversely, for a large enough number of fermions, long distance physics is governed by the appearance of an infrared fixed point, with its associated conformal invariance. This can be shown perturbatively, when $N_{f}$ is close to its upper bound insuring asymptotic freedom [23]. The conformal behavior is expected to persist over a finite range of $N_{f}>N_{f\,c}$, with the lower bound defining the edge of the conformal window. The determination of $N_{f\,c}$, and the characterization of the theory near it, is an interesting strong-coupling problem.
@@
Lattice field theory is the natural instrument to study the problem, and in recent years it has led to significant advances—see the review in Ref. [60] and references therein, as well as the more recent Refs. [31, 25, 59, 24, 19]. These studies explore a finite range of fermion masses that suppress infrared sensitivity to the lattice volume, and yield a discrete spectrum of bound states, as expected in confining theories. Some of these studies have reported evidence for the existence of a relatively light, flavor-singlet particle, together with a set of similarly light pseudoscalar particles, separated in mass from heavier bound states [7, 36, 43, 3, 16, 4, 56, 5, 54, 26, 6, 2, 42, 40, 41, 44, 45, 46, 21, 20, 22, 19].
@@
In a gauge theory outside of the conformal window, the pseudoscalars are naturally interpreted as pseudo-Nambu-Goldstone Bosons (pNGBs), associated with the spontaneous breaking of an approximate internal global symmetry acting on the matter fields. With the light flavor-singlet interpreted as a dilaton associated with the spontaneous breaking of approximate dilatation symmetry, the behavior of a gauge theory taken to be just below the edge of the window has been successfully described employing dilaton effective field theory (dEFT) [57, 49, 55, 53, 50, 8, 9, 29, 51, 30, 10, 48, 52]. See also the review in Ref. [13] and references therein, including the precursors in Refs. [58, 33], the early study in the context of electroweak symmetry breaking [47], as well as recent phenomenological applications in Refs. [11, 12, 28, 14, 27, 1, 64], and theoretical developments in Refs. [65, 38, 37, 34, 63, 62].
@@
In this paper, we expand on this idea, noting that the same dEFT can also be used just inside the conformal window, in the presence of a finite fermion mass. The dilaton and pNGBs continue to exist there as relatively light states. Therefore, dEFT can be employed to explore near-edge gauge theories without an inside-versus-outside bias, and hence used as a diagnostic tool to discriminate between confining and conformal behavior. We test this framework by applying it to two gauge theories in proximity to the lower edge of the conformal window. We start with the well studied $SU(3)$ gauge theory with $N_{f}=8$ Dirac fermions in the fundamental representation [26, 6], finding, as assumed in Refs. [57, 49, 55, 53, 50, 8, 9, 29, 51, 30, 10, 48, 52], that it is outside the conformal window. We then analyze the $SU(2)$ gauge theory coupled to $N_{f}=1$ Dirac fermion transforming in the adjoint representation [22, 19], finding evidence that it is inside.
@@
In Section II, we discuss the general framework. We then implement it to fit the lattice data for the above two gauge theories in Section III. In Section IV, we summarize our results and discuss future directions. We relegate to the Appendix some technical details about our fitting procedure, and tables reproducing the data we take from Refs. [26, 39] and [22, 18, 17].
@@ II Dilaton Effective Field Theory
The matrix field $\Sigma$ describes the pNGBs. The factor $\chi^{2}$ multiplying their kinetic term acts as a conformal compensator, and $P$ is a positive dimensionless parameter to be fit to lattice data. The final term describes explicit breaking of the internal global symmetry arising from the finite mass $m$ assigned to the fermions in the underlying gauge theory. The $\chi$ dependence of the terms in Eq. (II) follows from a spurion analysis, explained in Refs. [10, 13]. Heavier states and subleading interactions are neglected.
@@
We note in passing that our previous studies [10, 13] of the $SU(3)$ gauge theory with $N_{f}=8$ fermions in the fundamental representation were conducted with a constraint on the parameters of the potential $V(\chi)$ such that it has a symmetry breaking minimum, insuring that the spontaneous breaking of approximate conformal symmetry persists in the limit $R=0$. Here we proceed more generally, allowing lattice fits to determine whether or not $V(\chi)$ has a symmetry-breaking minimum, that is whether the theory is just outside or just inside the conformal window.
@@ III Fits to Lattice Data
Figure 3: Plot of $a^{3}M^{2}_{\pi}F_{\pi}$ versus $aF_{\pi}$, in the $SU(3)$ theory with $N_{f}=8$ fermions in the fundamental representation [26, 39]. The red curve is based on the central values of Table 1, with the five data points denoted by the black discs. The blue shaded region represents the uncertainty in the fitting curve, equivalent to the $1\sigma$ confidence level. The red curve reaches the horizontal axis at a nonzero value of $aF_{\pi}$, corresponding to the limit $m\rightarrow 0$ in a confining gauge theory. For larger values of $F_{\pi}\propto F_{d}$, we have $M_{\pi}^{2}F_{\pi}\propto V^{\prime}(F_{d})$. The curve is extrapolated downward, where it becomes negative, by continuing Eq. (10) to unphysical values of $F_{\pi}$, demonstrating that the region near the origin is unstable, as shown in the inset.
Table 1: Central values in the dEFT fit for the $SU(3)$ gauge theory with $N_{f}=8$ Dirac fermions in the fundamental representation. Lattice data for $M_{\pi}$, $M_{d}$, and $F_{\pi}$ is incorporated into this fit, for 5 different values of the underlying fermion mass $m$, corresponding to 15 data points. The lattice measurements are taken from Refs. [26, 39], and reproduced in Table 3. Dimensional quantities are presented in units of the lattice spacing, $a$.
@@ III.1 $SU(3)$ Gauge Theory with $N_{f}=8$ Fermions in the Fundamental Representation
The lattice data obtained from Refs. [26, 39] is summarized in Table 3 of Appendix A. The masses and decay constants of the $63$ pNGBs and the masses of the dilaton are listed there for $5$ values of the underlying fermion mass $m$. As noted in the Appendix, correlations are neglected. The choice of lattice coupling $\beta$ and the range of fermion mass $m$ employed in the lattice study lead to masses $M_{\pi}$ and $M_{d}$ well below the inverse lattice spacing, and also below the masses of heavier states, which are not included in the dEFT analysis.
Results for $A$, $B$, and $\Delta$ are shown in Figs. 1 and 2 in the form of $\Delta\chi^{2}$ contours. The favored upper left quadrants together correspond to values $\Delta<4$, $A>0$, and $B<0$. The central values of all $6$ fit parameters are shown in Table 1. Those for $y$ and $P$ are consistent with earlier studies. The central values of $\Delta$ (<4)<4), AA (positive), and $B$ (negative) imply that the theory lies just outside the conformal window, with $V(\chi)$ having a symmetry breaking minimum, as assumed in our earlier studies [13]. This is compatible with the expectation from the renormalization scheme invariant, multi-loop analysis of Ref. [61], as well as the analysis of Ref. [31], which combines lattice and perturbative inputs.
@@ III.2 $SU(2)$ Gauge Theory with $N_{f}=1$ Fermion in the Adjoint Representation
Figure 6: Plot of $a^{3}M^{2}_{\pi}F_{\pi}$ versus $aF_{\pi}\propto aF_{d}$, in the $SU(2)$ theory with $N_{f}=1$ fermion in the adjoint representation  [22, 18, 17], restricting attention to ensembles with $\beta=2.1$. The red curve is based on the central values of Table 2, using only the three lightest ensembles. The data points are depicted by the black discs. The red curve reaches the horizontal line at $F_{\pi}=0$, as shown in the inset, corresponding to the limit $m\rightarrow 0$ in an infrared conformal theory. Throughout the range, $a^{3}M^{2}_{\pi}F_{\pi}\propto V^{\prime}(F_{d})$. The blue shaded region represents the uncertainty in the fitting curve, equivalent to the $1\sigma$ confidence level.
Table 2: Central values for the dEFT fit for the $SU(2)$ gauge theory with $N_{f}=1$ Dirac fermion in the adjoint representation. The lattice data is taken from Refs. [22, 18, 17], and reproduced in Table 4, and has been restricted to the three ensembles with smallest fermion mass and $\beta=2.1$. Lattice data for $M_{\pi}$, $M_{d}$, $F_{\pi}$, and $m_{\rm PCAC}$ has been incorporated into this fit, corresponding to 9 data points. Dimensional quantities are presented in units of the lattice spacing, $a$.
@@
We again employ Eqs. (9)-(11) to fit the lattice data [22, 19, 18, 17]. In this case, with the spontaneous breaking pattern $SU(2)\rightarrow SO(2)$, there are $2$ pNGBs. We list the measurements performed with Wilson fermions and coupling $\beta=2.1$, for the largest available volumes, in Table 4 of Appendix A. We identify the fermion mass $m_{\rm PCAC}$, defined following Ref. [35], with the mass parameter $m$ in our fits. We include in our analysis only the three ensembles with the lightest mass, $aM_{\pi,d}<0.7$. We tested the stability of the fits by including one or more heavier ensembles. [^2] As discussed in Ref. [19], ensembles with $\beta>2.1$ yield a systematic decrease in the anomalous dimension of the fermion condensate, indicating to us that in those ensembles the Wilson fermion mass does not reach low enough values to assess the existence of an IR fixed point.
Results for $A$, $B$, and $\Delta$ are shown in Figs. 4 and 5 in the form of $\Delta\chi^{2}$ contours, with central values listed in Table 2. Among the central values, those for $y$ and $P$ are similar to those of Table 1. In particular, $y\approx 2$, as expected near the edge of the conformal window [32]. We again have $A>0$ and $B<0$, but now $\Delta>4$. Thus $V(\chi)$ turns down at large $\chi$, but the turn-down sets in well beyond the range of applicability of the dEFT.
@@ IV Summary and Discussion
We employed dilaton effective field theory (dEFT) as a diagnostic tool to discriminate between infrared-conformal and near-conformal, confining gauge theories. Our approach draws on spectroscopic lattice measurements performed with finite mass for the elementary fermions, in a range of parameter space in which both the pNGBs and the dilaton are clearly lighter than other states. This feature may be expected to emerge naturally in theories near the lower edge of the conformal window, for small enough fermion masses. We took the dEFT to apply smoothly through the transition between confinement and conformality.
We tested this approach with two lattice studies meeting these requirements. For the $SU(3)$ gauge theory coupled to $N_{f}=8$ Dirac fermions in the fundamental representation, we used measurements obtained with staggered fermions, restricted to the largest available lattice couplings and volumes [26, 6], to infer that the theory is outside the conformal window (near-conformal but confining in the $m\rightarrow 0$ limit). For the $SU(2)$ theory with $N_{f}=1$ fermions in the adjoint representation, we used data obtained with Wilson fermions [22, 19], by restricting attention to ensembles in which the relevant masses are significantly smaller than the lattice cutoff, to minimize lattice artifacts. Our analysis indicates that this theory is inside the conformal window (infrared conformal in the $m\rightarrow 0$ limit).
@@
Our results should be taken as preliminary, as indicated by the large uncertainty in some of the outputs—in particular the scaling dimension $\Delta$. Future precision lattice measurements, particularly for the $SU(2)$ theory, could overcome these limitations, by exploring lattice parameters with smaller fermion masses and less contamination from lattice artifacts. Our framework can be applied to other gauge theories. It would be interesting, for example, to apply our analysis to the $SU(3)$ theory coupled to $N_{f}=10$ Dirac fermions in the fundamental representation [56], as well as to lattice data for the theory with $N_{f}=2$ sextet fermions [42, 40, 41, 44, 45, 46], and to theories with other gauge groups, such as $Sp(4)$ [25, 24].